I have tried to work this equation, but get the wrong answer each time. HELP!
4-2x=9
thanks

Answer by Alan3354(69443)

I have tried to work this equation, but get the wrong answer each time. HELP!
4-2x=9
--------
Subtract 4
-2x = 5
Divide by -2
x = -5/2
What answer did you get? And how?